# Time Resolved Absorption of Six Chemical Species With MAROON-X Points to a Strong Drag in the Ultra Hot Jupiter TOI-1518 b

### v3

@@ Abstract
Context. Wind dynamics play a pivotal role in governing transport processes within planetary atmospheres, influencing atmospheric chemistry, cloud formation, and the overall energy budget. Understanding the strength and patterns of winds is crucial for comprehensive insights into the physics of ultra-hot Jupiter atmospheres. Current research has proposed different mechanisms that limit wind speeds in these atmospheres.
Aims. This study focuses on unraveling the wind dynamics and the chemical composition in the atmosphere of the ultra-hot Jupiter TOI-1518 b.
Methods. Two transit observations using the high-resolution ( 000), optical (spectral coverage between 490 and 920 nm) spectrograph MAROON-X were obtained and analyzed to explore the chemical composition and wind dynamics using the cross-correlation techniques, global circulation models, and atmospheric retrieval.
Results. We report the detection of 14 species in the atmosphere of TOI-1518 b through cross-correlation analysis. Vanadium Oxide was detected only with the new HyVO line list whereas Titanium Oxide was not detected. Additionally, we measure the time-varying cross-correlation trails for 6 different species, compare them with predictions from Global Circulation Models (GCM) and conclude that a strong drag is slowing the winds in TOI-1518 b’s atmosphere (). We find that the trails are species dependent. favors stronger drag than Fe, which we interpret as a sign of magnetic effects being responsible for the observed strong drag. Furthermore, we show that probes layers above the Roche lobe, leading to a qualitatively different trail than the other species. Finally, We use a retrieval analysis to further characterize the abundances of the different species detected. That analysis is refined thanks to the updated planetary mass of MJup we derived from the radial-velocity detection we also report here using SOPHIE data. We mesure an abundance of iron of corresponding to 0.07 to 1.62 solar enrichment. For the other elements, the retrievals appear to be biased, probably due to the different Kp/Vsys shifts between iron and the other elements, which we demonstrate for the case of .
@@ 1 Introduction
(Madhusudhan, 2019; Wordsworth & Kreidberg, 2022). Exoplanet atmosheres can be studied by observing their spectra either in emission (Chauvin et al., 2005; Swain et al., 2008), in transmission (Charbonneau et al., 2002) and soon in reflected light (Martins et al., 2013) from ground-based and space observatories. Recently, the sensitivity of
the JWST has begun to revolutionise this field, enabling extremely advanced studies of the fine structure and dynamics of the atmospheres of giant planets (Tsai et al., 2023; Coulombe et al., 2023). Thanks to their extended atmospheres, ultra-hot Jupiters (UHJs) are ideal targets for atmospheric characterization in transmission. These planets are very close to their stars and are tidally locked. This implies a significant day/night temperature gradient, which creates strong atmospheric circulation (Showman et al., 2020). Differences in temperature of several hundred degrees have been measured between the daysides and nightsides (Parmentier & Crossfield, 2018).
Due to their extreme temperature, volatile and refractory elements are accessible and detectable in such atmospheres. Indeed refractory species (with high condensation temperature) are expected to be gaseous in UHJs (Lothringer et al., 2018) when, in colder planets they are inaccessible because they condensed out of the gas phase. The measure of the refractory to volatile elemental ratio of these planets (e.g., O/Fe. C/Fe), recently emerged as a new powerful way to trace planet formation (Lothringer et al., 2021; Chachan et al., 2023; Pelletier et al., 2024; Smith et al., 2024).
@@
For the first time in 2010, high-resolution () spectroscopy was used to characterize the atmosphere of a transiting hot Jupiter by resolving individual molecular lines using CRIRES at VLT (Snellen et al., 2010). During the transit, the Doppler shifts caused by the planet’s rotation and the atmospheric winds allow lines formed in different parts of the planetary atmosphere to be spectroscopically separated (e.g. Nortmann et al. 2024). Recent ESPRESSO observations at VLT showed that the iron absorption lines of WASP-76b and WASP-121b, two canonical Ultra-Hot Jupiters (UHJs), are progressively blueshifted during the transit (Ehrenreich et al., 2020; Borsa et al., 2021).
@@ 2 Observations and Data reduction
We observed two transits of the UHJ TOI-1518 b with MAROON-X, a high-resolution () optical (spectral coverage between 490 and 920 nm) spectrograph at the 8.1-m Gemini-North observatory in Hawaii. Recent observations showed the capacity of MAROON-X to characterize UHJs by detecting ions and volatile and refractory elements. It has also allowed the description of the time variation of the atmospheric signal during transit (also known as the ”trail”) (Pelletier et al., 2023; Prinoth et al., 2023) and even to study some strong lines such as triplet (Prinoth et al., 2024). The observations were taken on 2022-08-13 and 2023-10-19 (program ID GN-2022B-Q-128 and GN-2023B-Q-127, PI: Parmentier). A summary of the observations is given in Table 1.
@@ 2.1 Data Quality
A principal component analysis (PCA) approach removes the telluric signal and residuals from the stellar correction. For most species we removed the first three principal components and verified that this was not significantly affecting our signal (see Fig. 24). This is not the case for the lines, which dominate the observed spectrum and are thus too affected by the PCA. For these we decided to not use any PCA correction.
We show in Fig. 2 two orders of MAROON-X red detectors before (top) and after (bottom) the division by the master out stellar spectrum. The strong stellar lines are efficiently removed, leaving apparent the planetary lines and the Doppler shadow effect (see sub-Section 3.4). In this wavelength range, the planetary lines are clearly visible, even without PCA corrections.
@@ 3.1 TOI-1518 system
TOI-1518 b is an ultra-hot Jupiter that was discovered by Cabot et al. (2021). It is a 1.875 R planet with an equilibrium temperature for zero albedo of 2546 K. The planet is misaligned, with an impact parameter of 0.9 (see Table.2), meaning that it doesn’t cross the zero-velocity planet of the star. The planet is precessing and the impact parameter has been shown to vary through time (Watanabe et al., 2024). Between our observations a change of 0.3 was expected, which is too small to affect our observations(from b = 0.91497 in 2019 to b= 0.8797 in 2022).The geometry of the system is presented in Fig. 18.
@@ 3.2 Template spectra for cross-correlation
The Cross-Correlation method is needed to detect the faint planetary lines in the residuals obtained after PCA. We cannot detect most of them directly except for a few individual lines (such as ). Fortunately, atoms and molecules have many spectral lines and a cross-correlation with a template boosts the signal, allowing a precise detection of the planetary spectrum. A template is needed to combine these lines. We generated synthetic spectra specific to TOI-1518 b using the parameters from Table. 2 and PetitRadtrans (Mollière et al., 2019), assuming a temperature of 2500K, with chemical abundances determined by equilibrium chemistry and calculated with FastChem (Stock et al., 2022).
@@
Alkaline metals and ions show individual strong lines, while other metals show line forests. Most of the signals except for the triplet are stronger in the blue part of the spectrum, which corresponds to the blue detector of MAROON-X. We decided to analyze each detector individually as if it were two different transits and then to sum every - map or CCF map.
@@ 3.4 Trails of the signal in CCF maps
Fig. 5 presents the time-variation of the cross-correlation function calculated with an iron template and shifted into the planetary rest frame. The dashed yellow line on the right underlines the expected position of the planet if the atmosphere is static and homogeneous. As the planet is misaligned, the Doppler shadow effect (already observed in Cont et al. 2021) only affects the planetary signal at the beginning of the transit and we have thus decided to not mask it nor to remove it.
@@
We further detect the time-varying trace of the CCF for six different species: Fe, , Ca, , Na, Mg (Fig. 7). All species show a qualitatively similar behavior to the Fe trace, with a blueshift over time.
@@ 3.5 Discussion
The detection of VO was made with a significance of 4.9 sigma using the newly released HyVO line list this year (Bowesman et al., 2024). VO is notoriously difficult to detect in exoplanet atmospheres; however, Pelletier et al. 2023 demonstrates the feasibility of such detections when employing a more accurate line list. For WASP-76b, the VOmyt line list was useful for the detection of . Fig. 8 illustrates that this line list was unsuccessful in retrieving VO signals in the observations of TOI-1518 b. This discrepancy can be attributed to a stronger signal in WASP-76b, along with the use of three transit observations compared to only two for TOI-1518 b. In this study, we demonstrate the superior capability of the HyVO line list, which successfully detected VO (see Fig. 8). The detected signal is blueshifted by as much as -12.11 0.22 km/s (see Table 3), which is significantly higher than the maximum blueshift observed for other species in TOI-1518 b, recorded at -5.16 0.57 km/s. Currently, it remains uncertain whether this difference is due to a shift in the line list itself or a physical shift in the atmosphere, potentially stemming from the varying localization of VO compared to metals and ions in the atmosphere of TOI-1518 b. The presence of strong optical absorbers such as VO and TiO in the atmosphere of UHJs is crucial for thermal inversion phenomena (Fortney et al., 2008). This finding underscores the significance of utilizing more accurate line lists for the detection of these molecules, especially in cases where the signals may be weaker than those observed in WASP-76b. Further studies employing this new line list could reveal the presence of VO in targets where it was previously undetectable, as highlighted in Borsa et al. (2021).
Detecting , , , and potentially raises questions. For example, is expected more in the hotter dayside atmosphere, whereas should be more present on the cooler limbs and nightside. However, the tracks of and seen in Figure 7 show a similar blueshifting trend, meaning that they are likely probing similar atmospheric regions.
@@
One of the main uncertainties of this work is the impact of the PCA step on the planetary signal. Fig. 24 shows the iron map for three components removed on the left, and on the right, it shows the maximum of the map in the function of the number of PCA components removed for three different boxes where measured the standard deviation of the map. The effect of PCA on high orbital velocity residuals is less than that on low velocity residuals, so calculating the standard deviation in the red box will underestimate the S/N. Conversely, the effect of PCA on low velocity residuals is much stronger, so the residuals will be smaller than those for high velocity, and the S/N will be overestimated. We then decided to calculate the standard deviation from the blue box, which mitigates this effect by taking the residuals at each velocity, but still far from the RM residuals or planetary signals, to avoid misinterpreting the standard deviation of the residual. The same method was used for each species observed.
@@
Several maps, such as Fe, Mg, Ca, or Cr, show a parasitic signal at km/s and km/s. This signal is a residual of the Rossiter-McLaughlin effect observed in the CCF map in Fig. 5. The Rossiter-McLaughlin effect presents an anti-correlation signal at negative . Therefore, it is not shown here, but the residual positive correlation visible in yellow around the anti-correlation signal in the CCF map is the parasite signal observed in the diagram mentioned previously. Possible biases due to Rossiter Mc-Laughing effect are present at phase below -5 degree and are represented in grey on the trail map of Fig. 7. The positive positive correlation at phases above 5 degree is far from the planetary signal 5 ( km/s). For the case of Mg, a signal is also visible at expected but = 100 km.s-1. This is due to a strong line in the Mg triplet. This is also visible in the CCF maps of both species, where the residual signal of the other can be observed as highlighted in Fig. 20.
@@
Fig. 22 highlights some limitations of the Gaussian profiles used to parameterize the iron trail of TOI-1518 b as most one-dimensional CCF do not follow a simple Gaussian profile. Then we decided to center the fitted Gaussian profile on the maximum of the 1D-CCF even if this resulted in a misestimation of the FWHM and probably of the error bars of the measured V. Another uncertainty might be due to the PCA applied to the data. Fig. 21 presents the iron trail of TOI-1518 b with different numbers of principal components removed. The square root of the variance of the Doppler shifts across all numbers of removed components, PCA has been added to the uncertainty quoted in the covariance matrix of the Gaussian fit obtained from scipy.optimize.curve_fit. The PCA does not change the transit trend even when many components are removed for all species except (Fig.23). The PCA is essential for detecting faint species. To maintain consistency, we removed the same number of components (three) from all species, even for those where it might not have been necessary. The only exception is , as the PCA significantly impacts the signal, even when fewer components are removed, due to its very high signal strength. Therefore, we decided not to apply PCA to in order to achieve a more robust analysis of its trail.
@@ 4 Comparison with global circulation models
To understand the physics behind the six time-resolved absorption trails we observed, we now compare our data with a suite of global circulation models.
@@ 4.1 Model description
These GCMs account for heat transport due to H2 dissociation and recombination (e.g., Bell & Cowan (2018); Komacek & Tan (2018);Tan & Komacek (2019); Roth et al. (2021)). H2 thermally dissociates on the dayside, after which atomic hydrogen gets advected to the nightside, where it recombines into H2 and releases latent heat. When the atmospheric circulation is predominantly eastward, most of this heat is pumped on the evening limb, resulting in a temperature asymmetry between the eastern and western regions of the atmosphere (first and second column in Fig.9). Drag restores energy to the atmosphere. Increasing drag strength (i.e lowering ) slows down winds in the atmosphere hindering this model’s heat transport. In strong drag cases, the temperature structure becomes symmetric, resulting in similar chemical compositions in the morning and evening limbs. (last column in Fig. 9). In the strong-drag model, there is only a day-to-night flow as the equatorial jet gets suppressed (Showman et al., 2013). Table 4 summarizes some other important parameters of the two SPARC/MITgcm models. We refer to Table 1 in Tan et al. (2024) for the full list of opacities considered in their radiative transfer (which include species such as and driving thermal inversion on the planet dayside). All models were run at a horizontal resolution of C32, corresponding to roughly 128 cells in longitude and 64 in latitude. Before computing phase-dependent spectra of the GCMs with gCMCRT, we bin the outputs down to 32 latitudes and 64 longitudes, as in Wardenier et al. (2021, 2023, 2024).

### v4

@@ Abstract
Context.
Aims. Wind dynamics play a pivotal role in governing transport processes within planetary atmospheres, influencing atmospheric chemistry, cloud formation, and the overall energy budget. Understanding the strength and patterns of winds is crucial for comprehensive insights into the physics of ultra-hot Jupiter atmospheres. Current research has proposed different mechanisms that limit wind speeds in these atmospheres.
Methods. This study focuses on unraveling the wind dynamics and the chemical composition in the atmosphere of the ultra-hot Jupiter TOI-1518 b.
Results. Two transit observations using the high-resolution ( 000), optical (spectral coverage between 490 and 920 nm) spectrograph MAROON-X were obtained and analyzed to explore the chemical composition and wind dynamics using the cross-correlation techniques, global circulation models (GCMs), and atmospheric retrieval.
Conclusions. We report the detection of 14 species in the atmosphere of TOI-1518 b through cross-correlation analysis. Vanadium Oxide was detected only with the new HyVO line list whereas Titanium Oxide was not detected. Additionally, we measure the time-varying cross-correlation trails for 6 different species, compare them with predictions from GCMs and conclude that a strong drag is slowing the winds in TOI-1518 b’s atmosphere (). We find that the trails are species dependent. Fe+ favors stronger drag than Fe, which we interpret as a sign of magnetic effects being responsible for the observed strong drag. Furthermore, we show that Ca+ probes layers above the Roche lobe, leading to a qualitatively different trail than the other species. Finally, We use a retrieval analysis to further characterize the abundances of the different species detected. Our analysis is refined thanks to the updated planetary mass of MJup we derived from new Sophie radial-velocity observations. We measure an abundance of iron of corresponding to 0.07 to 1.62 solar enrichment. For the other elements, the retrievals appear to be biased, probably due to the different Kp/Vsys shifts between iron and the other elements, which we demonstrate for the case of VO.
@@ 1 Introduction
(Madhusudhan, 2019; Wordsworth & Kreidberg, 2022). Exoplanet atmosheres can be studied by observing their spectra either in emission (Chauvin et al., 2005; Swain et al., 2008), in transmission (Charbonneau et al., 2002) and soon in reflected light (Martins et al., 2013) from ground-based and space observatories. Recently, the sensitivity of the JWST has begun to revolutionize this field, enabling extremely advanced studies of the fine structure and dynamics of the atmospheres of giant planets (Tsai et al., 2023; Coulombe et al., 2023). Thanks to their extended atmospheres, ultra-hot Jupiters (UHJs) are ideal targets for atmospheric characterization in transmission. These planets are very close to their stars and are tidally locked. This implies a significant day/night temperature gradient, which creates strong atmospheric circulation (Showman et al., 2020). Differences in temperature of several hundred degrees have been measured between the daysides and nightsides (Parmentier & Crossfield, 2018).
Due to their extreme temperature, volatile and refractory elements are accessible and detectable in such atmospheres. Indeed refractory species (with high condensation temperature) are expected to be gaseous in UHJs (Lothringer et al., 2018) when, in colder planets they are inaccessible because they condensed out of the gas phase. The measure of the refractory to volatile elemental ratio of these planets (e.g., O/Fe. C/Fe), recently emerged as a new powerful way to trace planet formation (Lothringer et al., 2021; Chachan et al., 2023; Pelletier et al., 2025; Smith et al., 2025).
@@
For the first time in 2010, high-resolution () spectroscopy was used to characterize the atmosphere of a transiting hot Jupiter by resolving individual molecular lines using CRIRES at VLT (Snellen et al., 2010). During the transit, the Doppler shifts caused by the planet’s rotation and the atmospheric winds allow lines formed in different parts of the planetary atmosphere to be spectroscopically separated (e.g. Nortmann et al. 2025). Recent ESPRESSO observations at VLT showed that the iron absorption lines of WASP-76b and WASP-121b, two canonical Ultra-Hot Jupiters (UHJs), are progressively blueshifted during the transit (Ehrenreich et al., 2020; Borsa et al., 2021).
@@ 2 Observations and data reduction
We observed two transits of the UHJ TOI-1518 b with MAROON-X, a high-resolution () optical (spectral coverage between 490 and 920 nm) spectrograph at the 8.1-m Gemini-North observatory in Hawaii. Recent observations showed the capacity of MAROON-X to characterize UHJs by detecting ions and volatile and refractory elements. It has also allowed the description of the time variation of the atmospheric signal during transit (also known as the ”trail”) (Pelletier et al., 2023; Prinoth et al., 2023) and even to study some strong lines such as Ca+ triplet (Prinoth et al., 2024). The observations were taken on 2022-08-13 and 2023-10-19 (program ID GN-2022B-Q-128 and GN-2023B-Q-127, PI: Parmentier). A summary of the observations is given in Table 1.
@@
A principal component analysis (PCA) approach removes the telluric signal and residuals from the stellar correction. For most species we removed the first three principal components and verified that this was not significantly affecting our signal (see Fig. 24). This is not the case for the Ca+ lines, which dominate the observed spectrum and are thus too affected by the PCA. For these we decided to not use any PCA correction.
We show in Fig. 2 two orders of MAROON-X red detectors before (top) and after (bottom) the division by the master out stellar spectrum. The strong stellar Ca+ lines are efficiently removed, leaving apparent the planetary lines and the Doppler shadow effect (see sub-Section 3.4). In this wavelength range, the planetary Ca+ lines are clearly visible, even without PCA corrections.
@@ 3.1 TOI-1518 system
TOI-1518 b is an ultra-hot Jupiter that was discovered by Cabot et al. (2021). It is a 1.875 R planet with an equilibrium temperature for zero albedo of 2546 K. The planet is misaligned, with an impact parameter of 0.9 (see Table.2), meaning that it doesn’t cross the zero-velocity plane of the star. The planet is precessing and the impact parameter has been shown to vary through time (Watanabe et al., 2024). Between our observations a change of 0.03 was expected, which is too small to affect our observations(from b = 0.91497 in 2019 to b= 0.8797 in 2022).The geometry of the system is presented in Fig. 18.
@@ 3.2 Template spectra for cross-correlation
The Cross-Correlation method is needed to detect the faint planetary lines in the residuals obtained after PCA. We cannot detect most of them directly except for a few individual lines (such as Ca+). Fortunately, atoms and molecules have many spectral lines and a cross-correlation with a template boosts the signal, allowing a precise detection of the planetary spectrum. A template is needed to combine these lines. We generated synthetic spectra specific to TOI-1518 b using the parameters from Table. 2 and PetitRadtrans (Mollière et al., 2019), assuming a temperature of 2500K, with chemical abundances determined by equilibrium chemistry and calculated with FastChem (Stock et al., 2022).
@@
Alkaline metals and ions show individual strong lines, while other metals show line forests. Most of the signals except for the Ca+ triplet are stronger in the blue part of the spectrum, which corresponds to the blue detector of MAROON-X. We decided to analyze each detector individually as if it were two different transits and then to sum every - map or Cross-Correlation Function (CCF) map.
@@ 3.3 Species detected in the - maps
Notes : Amplitude (Amp) corresponds to the best-fit line depth of the absorbing species above the spectral continuum. The amplitude is expressed in units of signal-to-noise ratio (), as each map is divided by the standard deviation as explained in Sect. 3.2. Detections range from 4 to 21.1 . corresponds to the radial velocity of the line center, as measured in the rest frame of the stellar system (same as ). FWHM denotes the Gaussian Full-Width at Half-Maximum. The other parameters are expressed in km/s. The value of is calculated by performing a Gaussian fit at the best fit . The error bars are determined from the covariance matrix of the Gaussian fit.
@@ 3.4 Trails of the signal in CCF maps
Fig. 5 presents the time-variation of the CCF calculated with an iron template and shifted into the planetary rest frame. The dashed yellow line on the right underlines the expected position of the planet if the atmosphere is static and homogeneous. As the planet is misaligned, the Doppler shadow effect (already observed in Cont et al. 2021) only affects the planetary signal at the beginning of the transit and we have thus decided to not mask it nor to remove it.
@@
We further detect the time-varying trace of the CCF for six different species: Fe, Fe+, Ca, Ca+, Na, Mg (Fig. 7). All species show a qualitatively similar behavior to the Fe trace, with a blueshift over time.
@@ 3.5 Discussion
The detection of VO was made with a significance of 4.9 sigma using the newly released HyVO line list this year (Bowesman et al., 2024). VO is notoriously difficult to detect in exoplanet atmospheres; however, Pelletier et al. 2023 demonstrates the feasibility of such detections when employing a more accurate line list. For WASP-76b, the VOmyt line list was useful for the detection of VO. Fig. 8 illustrates that this line list was unsuccessful in retrieving VO signals in the observations of TOI-1518 b. This discrepancy can be attributed to a stronger VO signal in WASP-76b, along with the use of three transit observations compared to only two for TOI-1518 b. In this study, we demonstrate the superior capability of the HyVO line list, which successfully detected VO (see Fig. 8). The detected signal is blueshifted by as much as -12.11 0.22 km/s (see Table 3), which is significantly higher than the maximum blueshift observed for other species in TOI-1518 b, recorded at -5.16 0.57 km/s. Currently, it remains uncertain whether this difference is due to a shift in the line list itself or a physical shift in the atmosphere, potentially stemming from the varying localization of VO compared to metals and ions in the atmosphere of TOI-1518 b. The presence of strong optical absorbers such as VO and TiO in the atmosphere of UHJs is crucial for thermal inversion phenomena (Fortney et al., 2008). This finding underscores the significance of utilizing more accurate line lists for the detection of these molecules, especially in cases where the signals may be weaker than those observed in WASP-76b. Further studies employing this new line list could reveal the presence of VO in targets where it was previously undetectable, as highlighted in Borsa et al. (2021).
Detecting Fe/Fe+, Ca/Ca+, Ti/Ti+, and potentially V/VO raises questions. For example, Fe+ is expected more in the hotter dayside atmosphere, whereas Fe should be more present on the cooler limbs and nightside. However, the tracks of Fe and Fe+ seen in Fig. 7 show a similar blueshifting trend, meaning that they are likely probing similar atmospheric regions.
@@
One of the main uncertainties of this work is the impact of the PCA step on the planetary signal. Fig. 24 shows the iron Kp-Vres map for three components removed on the left, and on the right, it shows the maximum of the Kp-Vres map in the function of the number of PCA components removed for three different boxes where measured the standard deviation of the map. The effect of PCA on high orbital velocity residuals is less than that on low velocity residuals, so calculating the standard deviation in the red box will underestimate the S/N. Conversely, the effect of PCA on low velocity residuals is much stronger, so the residuals will be smaller than those for high velocity, and the S/N will be overestimated. We then decided to calculate the standard deviation from the blue box, which mitigates this effect by taking the residuals at each velocity, but still far from the RM residuals or planetary signals, to avoid misinterpreting the standard deviation of the residual. The same method was used for each species observed.
@@
Several Kp-Vres maps, such as Fe, Mg, Ca, or Cr, show a parasitic signal at K km/s and V km/s. This signal is a residual of the Rossiter-McLaughlin effect observed in the CCF map in Fig. 5. The Rossiter-McLaughlin effect presents an anti-correlation signal at negative . Therefore, it is not shown here, but the residual positive correlation visible in yellow around the anti-correlation signal in the CCF map is the parasite signal observed in the Kp-Vres diagram mentioned previously. Possible biases due to Rossiter Mc-Laughing effect are present at phase below -5 degree and are represented in grey on the trail map of Fig. 7. The positive correlation at phases above 5 degree is far from the planetary signal 5 ( km/s). For the case of Mg, a signal is also visible at expected Kp but Vres = 100 km.s-1. This is due to a strong Fe+ line in the Mg triplet. This is also visible in the CCF maps of both species, where the residual signal of the other can be observed as highlighted in Fig. 20.
@@
Fig. 22 highlights some limitations of the Gaussian profiles used to parameterize the iron trail of TOI-1518 b as most one-dimensional CCF do not follow a simple Gaussian profile. Then we decided to center the fitted Gaussian profile on the maximum of the 1D-CCF even if this resulted in a misestimation of the FWHM and probably of the error bars of the measured V. Another uncertainty might be due to the PCA applied to the data. Fig. 21 presents the iron trail of TOI-1518 b with different numbers of principal components removed. The square root of the variance of the Doppler shifts across all numbers of removed components, PCA has been added to the uncertainty quoted in the covariance matrix of the Gaussian fit obtained from scipy.optimize.curve_fit. The PCA does not change the transit trend even when many components are removed for all species except Ca+ (Fig.23). The PCA is essential for detecting faint species. To maintain consistency, we removed the same number of components (three) from all species, even for those where it might not have been necessary. The only exception is Ca+, as the PCA significantly impacts the signal, even when fewer components are removed, due to its very high signal strength. Therefore, we decided not to apply PCA to Ca+ in order to achieve a more robust analysis of its trail.
@@ 4 Comparison with global circulation models
To understand the physics behind the six time-resolved absorption trails we observed, we now compare our data with a suite of GCMs.
@@ 4.1 Model description
These GCMs account for heat transport due to H2 dissociation and recombination (e.g., Bell & Cowan (2018); Komacek & Tan (2018);Tan & Komacek (2019); Roth et al. (2021)). H2 thermally dissociates on the dayside, after which atomic hydrogen gets advected to the nightside, where it recombines into H2 and releases latent heat. When the atmospheric circulation is predominantly eastward, most of this heat is pumped on the evening limb, resulting in a temperature asymmetry between the eastern and western regions of the atmosphere (first and second column in Fig.9). Drag restores energy to the atmosphere. Increasing drag strength (i.e lowering ) slows down winds in the atmosphere hindering this model’s heat transport. In strong drag cases, the temperature structure becomes symmetric, resulting in similar chemical compositions in the morning and evening limbs. (last column in Fig. 9). In the strong-drag model, there is only a day-to-night flow as the equatorial jet gets suppressed (Showman et al., 2013). Table 4 summarizes some other important parameters of the two SPARC/MITgcm models. We refer to Table 1 in Tan et al. (2024) for the full list of opacities considered in their radiative transfer (which include species such as TiO and Fe driving thermal inversion on the planet dayside). All models were run at a horizontal resolution of C32, corresponding to roughly 128 cells in longitude and 64 in latitude. Before computing phase-dependent spectra of the GCMs with gCMCRT, we bin the outputs down to 32 latitudes and 64 longitudes, as in Wardenier et al. (2021, 2023, 2024).